# Matching with Generalized Sequential Reserves

### v2

@@ 5.1. Affirmative Action in India
Horizontal reservations are provided to other marginalized groups, such as women and individuals with disabilities [^7] For an in-depth analysis, see Sönmez and Yenmez (2022).. The SCI’s judgment in Indra Sawhney (1992) mandates that horizontal reservations must be implemented within each vertical category on a minimum guarantee basis. That is, in each vertical category, horizontally reserved positions must be allocated before any unreserved positions. However, the SCI judgments have left it ambiguous whether an individual eligible for multiple types of horizontal reservation is counted against a single horizontal category or multiple categories concurrently.
Let $H=\{h_{1},\ldots,h_{L}\}$ be the set of horizontal types. The correspondence $\rho:\>I\rightrightarrows H\cup\{\emptyset\}$ represents the horizontal types of individuals. $\rho(i)\subseteq H\cup\{\emptyset\}$ denotes the set of horizontal types that the individual $i$ can claim. $\rho(i)=\emptyset$ means that $i$ has no horizontal type. We denote by $\kappa_{v}^{j}$ the number of positions reserved for horizontal type $h_{j}\in H$ in the vertical category $v\in V$. The vector $\kappa_{v}^{s}\equiv(\kappa_{v}^{1},...,\kappa_{v}^{L})$ denotes the horizontal reservations in the vertical category $v\in V$ in the institution $s$. Let $\kappa_{s}\equiv\left\{\kappa_{v}^{s}\right\}_{v\in V}$ denote the horizontal reservations of the institution $s$.
@@
SC and ST are classified as hard reserves, which means that the vacant positions reserved for these groups cannot be transferred to other categories. Although reservations for OBC are treated as hard reserves in the context of governmental job allocations, they were suggested and have been implemented as soft reserves in admissions to public colleges, pursuant to the landmark SCI decision in Ashoka Kumar Thakur vs. Union of India (2008). [^8] The judgment is available at https://indiankanoon.org/doc/1219385/. The SCI rulings did not provide an explicit procedural framework for de-reserving unfilled OBC positions. The lack of clear rules has led to various inconsistent practices in school admissions in India (Aygün and Turhan (2022)).
@@
Each institution $s$ has a strict priority order $\succ_{s}$ over ${}\mathcal{I}\cup\{\emptyset\}$. We write $i\succ_{s}j$ to mean that the applicant $i$ has a higher priority than $j$ at $s$. Similarly, we write $i\succ_{s}\emptyset$ to say that applicant $i$ is acceptable for $s$. $\emptyset\succ_{s}i$ means that the applicant $i$ is unacceptable for $s$. The profile of institutions’ priorities is denoted $\succ=(\succ_{s_{1}},...,\succ_{s_{m}})$. For each institution $s\in\mathcal{S}$, the merit ordering for individuals of type $r\in\mathcal{R}$, denoted by $\succ_{s}^{r}$, is obtained from $\succ_{s}$ in a straightforward manner as follows:
  • for $i,j\in\mathcal{I}$ such that $t_{i}=r$, $t_{j}\neq r$, $i\succ_{s}\emptyset$, and $j\succ_{s}\emptyset$, we have $i\succ_{s}^{r}\emptyset\succ_{s}^{r}j$, where $\emptyset\succ_{t}^{r}j$ means individual $j$ is unacceptable for category $r$ at institution $s$.
@@
We define $\mathcal{X}\equiv\underset{t(i)=g}{\bigcup}\{i\}\times S\times\{o\}\underset{t(i)\in R}{\bigcup}\{i\}\times S\times\{r,o\}$ as the set of all contracts. Each contract $x\in\mathcal{X}$ is between an individual $\mathbf{i}(x)$ and an institution $\mathbf{s}(x)$ and specifies a vertical category $\mathbf{t}(x)$ in which the individual $\mathbf{i}(x)$ is admitted.
Each individual in category g has a preference for $\mathcal{S\times}\left\{o\right\}\cup\left\{\emptyset\right\}$. Each member of the reserve category $r$ has a preference over $\mathcal{S}\times\left\{r,o\right\}\cup\left\{\emptyset\right\}$, where $\emptyset$ denotes the outside option. [^9] We assume that individuals do not have preferences for horizontal types as there is no evidence that individuals care about the horizontal types under which they are admitted. All discussion on reservation policy in India centers on vertical reservations. We write $(s,v)P_{i}(s^{{}^{\prime}},v^{{}^{\prime}})$ to mean that the individual $i$ strictly prefers admission to institution $s$ through the vertical category $v$ to admission to institution $s^{{}^{\prime}}$ via vertical category $v^{{}^{\prime}}$. The at-least-as-well relation $R_{i}$ is obtained from $P_{i}$ as follows: $(s,v)R_{i}(s^{{}^{\prime}},v^{{}^{\prime}})$ if and only if $(s,v)P_{i}(s^{{}^{\prime}},v^{{}^{\prime}})$ or $(s,v)=(s^{{}^{\prime}},v^{{}^{\prime}})$. An institution and vertical category pair $(s,v)$ is acceptable to individual $i$ if it is at least as good as the outside option $\emptyset$ and is unacceptable to her if it is worse than the outside option $\emptyset$.
@@
We consider one-to-one horizontal matching, where an admitted applicant with multiple eligible horizontal types is counted against only one of these types. Sönmez and Yenmez (2022) proposed the meritorious horizontal choice rule, denoted $C^{m}$, in this setting. Within each vertical category $v$, the authors consider the assignment of horizontally reserved positions $H^{v}=\bigcup_{j=1}^{L}H_{j}^{v}$ to a set of individuals $I$ as a one-to-one matching problem, where $H_{j}^{v}$ denotes the set of positions reserved for horizontal type $h_{j}\in H$ in vertical category $v$. The number of positions in $H_{j}^{v}$ is denoted by $\kappa_{v}^{j}$.
Given a vertical category $v\in\mathcal{V}$ and a set of individuals $I$ such that $t_{i}=v$ for all $i\in I$, define the one-to-one matching between the individuals and horizontally reserved positions as a mapping $\mu:I\rightarrow H^{v}\bigcup\left\{\emptyset\right\}$ such that for any $i\in I$ and $h_{j}\in H$,
@@
Given a vertical category $v$, a set of applicants $A$ such that $\mathbf{v}(i)=v$ for all $i\in A$, a profile of the horizontal types of applicants $\left(\rho(i)\right)_{i\in A}$, the meritorious horizontal choice rule $C^{m}$ selects the applicants as follows:
@@
Fleiner (2001) showed that the first stage of $C^{m}$ satisfies the substitute property. Sönmez and Yenmez (2022) proved that $C^{m}$ satisfies size monotonicity. We prove that $C^{m}$ also satisfies the quota monotonicity. Thus, any aggregate choice rule in which the divisions select according to $C^{m}$ under a monotone transfer policy is in the GS family.
@@
According to the aggregate choice rule with transfer, $C^{mT}$, vacant OBC positions are reverted to an open category and filled following merit scores. The rule $C^{mT}$ is an GS choice rule and can be written as
$C_{s}^{mT}\equiv\left(\left(C_{v}^{m}\right)_{v\in\mathcal{V}},q_{s}\right),$
@@
  • the precedence sequence is $o-SC-ST-OBC-EWS-D$, where $D$ denotes “de-reserved positions” that are reverted from OBC, and
  • $q_{s}=\left(q_{s}^{o},q_{s}^{SC},q_{s}^{ST},q_{s}^{OBC},q_{s}^{EWS},q_{s}^{D}\right)$ such that $q_{s}^{v}=\overline{q}_{s}^{v}$ for all $v\in\left\{o,SC,ST,OBC,EWS\right\}$, and $q_{s}^{D}=r_{OBC}$ is the number of vacant OBC positions.
When each vertical category $v\in\mathcal{V}$ chooses applicants through $C_{v}^{m}$, under any monotonic transfer policy, the aggregate choice rules of the institutions are in the GS family.
@@
$C_{s}^{mT}$ belongs to the GS family.
[…]
@@ 5.2. Affirmative Action in Chinese High School Admissions
In China, public high schools are subject to an affirmative action policy in which high schools reserve a proportion of their available seats for graduates of low-performing middle schools. High school admissions are merit-based and highly competitive. The affirmative action policy requires high schools to reserve 50% or more of their seats exogenously to middle schools.
Hu and Zhang (2025) study a novel aspect of the Chinese affirmative action policy, called the “verifiability," that has been implemented in almost every city in China. Verifiability of an allocation mechanism ensures that students can verify their assignments using only their private information (e.g., their preferences over schools) and publicly observed cutoff scores. It is an important aspect of the transparency and accountability of the admission process. The authors analyze several choice rules that are used in Chinese cities. To describe them, we first introduce their notation.
There is a finite set of students $\mathcal{I}$. Let $q_{s}$ be the capacity of the high school $s$. Let $\mathcal{M}$ be the finite set of middle schools. The high school $s$ reserves $q_{s}^{m}$ seats for students from the middle school $m$. We assume that $\sum_{m\in M}q_{s}^{m}<q_{s}$, where $q_{s}^{o}=q_{s}-\sum_{m\in M}q_{s}^{m}$ denotes the number of open seats in high school $s$. The function $\tau:\mathcal{I}\rightarrow\mathcal{M}$ is such that $\tau(i)$ denotes the middle school student $i$ belongs. Let $\mathcal{I}^{m}=\tau^{-1}(m)$ and $I^{m}\equiv I\cap\mathcal{I}^{m}$.
There are two deferred acceptance-based mechanisms that allocate open and reserved seats simultaneously. These mechanisms use the choice rules $\mathcal{C}^{SimRO}$ and $\mathcal{C}^{SimORO}$ we define below.
5.2.1. The Simultaneous-reserve-open Choice Rule $\mathcal{C}^{SimRO}$
This choice rule is used in cities such as Guiyan, Qingdao, and Tianjin for high school admissions in China.
For any given set $I\subseteq\mathcal{I}$ of students, the choice rule $\mathcal{C}^{SimRO}$ selects the students as follows:
@@ 5.2.1. The Simultaneous-reserve-open Choice Rule $\mathcal{C}^{SimRO}$
Note that unfilled reserved seats are allocated as open seats in the second stage. The choice rule $\mathcal{C}^{SimRO}$ is in the GS family.
5.2.2. The Simultaneous-open-reserve-open Choice Rule $\mathcal{C}^{SimORO}$
In Chinese cities Taiyuan, Zengzhou, and Hefei, among many others, for high school admissions, the deferred acceptance mechanism with the choice rule $\mathcal{C}^{SimORO}$ is used.
For any given set $I\subseteq\mathcal{I}$ of students, the choice rule $\mathcal{C}^{SimORO}$ selects the students in three steps.
@@ 5.2.2. The Simultaneous-open-reserve-open Choice Rule $\mathcal{C}^{SimORO}$
The unfilled reserved seats in the second stage are allocated as open seats in the third stage. The choice rule $\mathcal{C}^{SimORO}$ is in the GS family.
Deferred Acceptance Induced Choice Rules
Another version of the deferred acceptance mechanism used for high school admissions in some Chinese cities differs from the two mechanisms above in that each school is divided into two subschools, open subschool and reserved subschools. For school $s\in\mathcal{S}$, the open subschool $s^{o}$ has capacity $q_{S}^{o}$, and the reserved subschool $S^{r}_{s}$ has the capacity vector $q_{s}^{r}=(q_{S}^{m})_{m\in\mathcal{M}}$. We instead model it in matching with contracts framework.
Students are required to submit a preference list of subschools. Then, the student-proposing DA coupled with the following choice rules is run to find a matching. In each round of the DA, each unassigned student applies to the highest-ranked subschool in their preference lists.
@@
Consider only the contracts in $X_{r}$. For each $m\in\mathcal{M}$, choose contracts of the highest-ranked students in $(X_{r})_{i}\cap I^{m}$ up to capacity $q_{s}^{m}$.
This rule does not allocate reserved positions that are left unfilled. Thus, it is wasteful. Note that $\mathcal{C}^{SimSep}$ is in the GS family.
@@
Consider only the contracts in $X_{r}$. For each $m\in\mathcal{M}$, choose contracts of the highest-ranked students in $(X_{r})_{i}\cap I^{m}$ up to capacity $q_{s}^{m}$.
@@
$\mathcal{C}^{SimFlex}$ is a GS choice rule.
@@
Let $I$ be the set of students and $s$ be the only school. Define $X=I\times\{s\}\times\{o,r\}$. Assume that each student $i\in I$ ranks the contract with the open seat over the contract with the reserved seat. Run the student proposing DA while the school $s$ selects students using $\mathcal{C}^{SimFlex}$. The result of the DA procedure defines $\mathcal{C}^{SimFlex}(I)$.
Hu and Zhang (2025) show that $\mathcal{C}^{SimOR}$ does not belong to the class of slot-specific priorities choice rules of Kominers and Sönmez (2016). On the other hand, the GS family contains $\mathcal{C}^{SimOR}$ naturally because $\mathcal{C}^{SimOR}$ satisfies the substitutes condition and is $q$-acceptant, which implies size monotonicity. [^10] Since $\mathcal{C}^{SimOR}$ satisfies the substitutes condition and size monotonicity, it can be considered as a single division’s choice rule.

### v3

@@ 5.1. Affirmative Action in India
Horizontal reservations are provided to other marginalized groups, such as women and individuals with disabilities [^12] For an in-depth analysis, see sonmez2022affirmative.. The SCI’s judgment in Indra Sawhney (1992) mandates that horizontal reservations must be implemented within each vertical category on a minimum guarantee basis. That is, in each vertical category, horizontally reserved positions must be allocated before any unreserved positions are allocated. However, the SCI judgments have left it ambiguous whether an individual eligible for multiple types of horizontal reservation is counted against a single horizontal category or multiple categories concurrently.
Let $H=\{h_{1},\ldots,h_{L}\}$ be the set of horizontal types. The correspondence $\rho:\>I\rightrightarrows H\cup\{G\}$ represents the horizontal types of individuals. $\rho(i)\subseteq H\cup\{G\}$ denotes the set of horizontal types that the individual $i$ can claim. $\rho(i)=G$ means that $i$ has no horizontal type. We denote by $\kappa_{v}^{j}\in\mathbb{Z}_{\geq 0}$ the number of positions reserved for horizontal type $h_{j}\in H$ in the vertical category $v\in V$. The vector $\kappa_{v}^{s}\equiv(\kappa_{v}^{1},...,\kappa_{v}^{L})\in\mathbb{Z}_{\geq 0}^{L}$ denotes the horizontal reservations in the vertical category $v\in V$ in the institution $s$. Let $\Gamma{s}\equiv\left\{\kappa_{v}^{s}\right\}_{v\in V}\in\mathbb{Z}_{\geq 0}^{L\times|V|}$ denote the horizontal reservations of the institution $s$.
@@
SC and ST are classified as hard reserves, which means that vacant positions reserved for these groups cannot be transferred to other categories. Although reservations for OBC are treated as hard reserves in the context of governmental job allocations, they were suggested and have been implemented as soft reserves in admissions to public colleges, pursuant to the landmark SCI decision in Ashoka Kumar Thakur vs. Union of India (2008). [^13] The judgment is available at https://indiankanoon.org/doc/1219385/. The SCI rulings did not provide an explicit procedural framework for de-reserving unfilled OBC positions. The lack of clear rules has led to various inconsistent practices in school admissions in India (aygun2022dereserve).
@@
Each institution $s$ has a strict priority order $\succ_{s}$ over $\mathcal{I}\cup\{\emptyset_{s}\}$. We write $i\succ_{s}j$ to mean that the applicant $i$ has a higher priority than $j$ at $s$. Similarly, we write $i\succ_{s}\emptyset_{s}$ to say that applicant $i$ is acceptable for $s$. $\emptyset_{s}\succ_{s}i$ means that the applicant $i$ is unacceptable for $s$. The profile of institutions’ priorities is denoted $\succ=(\succ_{s_{1}},...,\succ_{s_{m}})$. For each institution $s\in\mathcal{S}$, the merit ordering for individuals of type $r\in\mathcal{R}$, denoted by $\succ_{s}^{r}$, is obtained from $\succ_{s}$ in a straightforward manner as follows:
  • for $i,j\in\mathcal{I}$ such that $t_{i}=r$, $t_{j}\neq r$, $i\succ_{s}\emptyset_{s}$, and $j\succ_{s}\emptyset_{s}$, we have $i\succ_{s}^{r}\emptyset_{s}\succ_{s}^{r}j$, where $\emptyset_{s}\succ_{t}^{r}j$ means individual $j$ is unacceptable for category $r$ at institution $s$.
@@
We define $\mathcal{X}\equiv\underset{t(i)=g}{\bigcup}\{i\}\times S\times\{o\}\underset{t(i)\in R}{\bigcup}\{i\}\times S\times\{t(i),o\}$ as the set of all contracts. Each contract $x\in\mathcal{X}$ is between an individual $\mathbf{i}(x)$ and an institution $\mathbf{s}(x)$ and specifies a vertical category $\mathbf{t}(x)$ in which the individual $\mathbf{i}(x)$ is admitted.
Each individual $i$ with $t(i)=g$ has a preference for $\mathcal{S\times}\left\{o\right\}\cup\left\{\emptyset_{i}\right\}$. Each individual $i$ with $t(i)=r\in R$ has a preference over $\mathcal{S}\times\left\{r,o\right\}\cup\left\{\emptyset_{i}\right\}$, where $\emptyset_{i}$ denotes the outside option for individual ii. [^14] We assume that individuals do not have preferences for horizontal types as there is no evidence that individuals care about the horizontal types under which they are admitted. All discussion on reservation policy in India centers on vertical reservations. We write $(s,v)P_{i}(s^{{}^{\prime}},v^{{}^{\prime}})$ to mean that the individual $i$ strictly prefers admission to institution $s$ through the vertical category $v$ to admission to institution $s^{{}^{\prime}}$ via vertical category $v^{{}^{\prime}}$. The at-least-as-well relation $R_{i}$ is obtained from $P_{i}$ as follows: $(s,v)R_{i}(s^{{}^{\prime}},v^{{}^{\prime}})$ if and only if $(s,v)P_{i}(s^{{}^{\prime}},v^{{}^{\prime}})$ or $(s,v)=(s^{{}^{\prime}},v^{{}^{\prime}})$. An institution and vertical category pair $(s,v)$ is acceptable to individual $i$ if it is at least as good as the outside option $\emptyset_{i}$ and is unacceptable to her if it is worse than the outside option $\emptyset_{i}$.
@@
We consider one-to-one horizontal matching, where an admitted applicant with multiple eligible horizontal types is counted against only one of these types. sonmez2022affirmative proposed the meritorious horizontal choice rule, denoted $C^{m}$, in this setting. Within each vertical category $v$, the authors consider the assignment of horizontally reserved positions $H^{v}=\bigcup_{j=1}^{L}H_{j}^{v}$ to a set of individuals $I$ as a one-to-one matching problem, where $H_{j}^{v}$ denotes the set of positions reserved for horizontal type $h_{j}\in H$ in vertical category $v$. The number of positions in $H_{j}^{v}$ is denoted by $\kappa_{v}^{j}$.
Given a vertical category $v\in\mathcal{V}$ and a set of individuals $I$ such that $t(i)=v$ for all $i\in I$, define the one-to-one matching between the individuals and horizontally reserved positions as a mapping $\mu:I\rightarrow H^{v}\bigcup\left\{\emptyset\right\}$ such that for any $i\in I$ and $h_{j}\in H$,
@@
Given a vertical category $v$, a set of applicants $A$ such that $t(i)=v$ for all $i\in A$, a profile of the horizontal types of applicants $\left(\rho(i)\right)_{i\in A}$, the meritorious horizontal choice rule $C^{m}$ selects the applicants as follows:
@@
fleiner2001matroid showed that the first stage of $C^{m}$ satisfies the substitute property. sonmez2022affirmative proved that $C^{m}$ satisfies size monotonicity. We prove that $C^{m}$ also satisfies the quota monotonicity. Thus, any aggregate choice rule in which the divisions select according to $C^{m}$ under a monotone transfer policy is in the GSq family.
@@
According to the aggregate choice rule with transfer, $C^{mT}$, vacant OBC positions are reverted to an open category and filled following merit scores. The rule $C^{mT}$ can be written as
$C_{s}^{mT}\equiv\left((\left(C_{v}^{m}\right)_{v\in\mathcal{V}},C_{D}^{m}),q_{s}\right),$
@@
  • the precedence sequence is oSCSTOBCEWSDo-SC-ST-OBC-EWS-D, where $D$ denotes “de-reserved positions” that are reverted from OBC, and
  • $q_{s}=\left(q_{s}^{o},q_{s}^{SC},q_{s}^{ST},q_{s}^{OBC},q_{s}^{EWS},q_{s}^{D}\right)$ such that $q_{s}^{v}=\overline{q}_{s}^{v}$ for all $v\in\left\{o,SC,ST,OBC,EWS\right\}$, and $q_{s}^{D}=r_{OBC}$ is the number of vacant OBC positions, and
  • $C_{C}^{m}$ is a $q_{s}^{D}$-responsive choice rule induced by the merit ranking of individuals.
Our next result states that when each vertical category $v\in\mathcal{V}$ chooses applicants through $C_{v}^{m}$, under any monotonic transfer policy, the aggregate choice rule $C_{s}^{mT}$ is in the GSq family.
@@
$C_{s}^{mT}$ belongs to the GSq family.
An immediate corollary of Proposition 1 is as follows.
Corollary 2.
Let $\Phi^{mT}$ be the COM under the profile of choice rules $C^{mT}=\left(C_{s}^{mT}\right)_{s\in\mathcal{S}}$. Then $\Phi^{mT}$ is stable and strategy-proof for individuals.
@@ 5.2. Affirmative Action in Chinese High School Admissions
In China, public high schools are subject to an affirmative action policy in which high schools reserve a proportion of their available seats for graduates of low-performing middle schools. High school admissions are merit-based and highly competitive. The affirmative action policy requires high schools to reserve 50% or more of their seats exogenously to middle schools. hu2025verifiable study a novel aspect of the Chinese affirmative action policy, called “verifiability," which has been implemented in almost every city in China. [^15] Verifiability of an allocation mechanism ensures that students can verify their assignments using only their private information (e.g., their preferences over schools) and publicly observed cutoff scores; It is an important aspect of the transparency and accountability of the admission process. The authors analyze several choice rules that are used in Chinese cities. They also propose new choice rules. To define them, we first introduce their notation.
There is a finite set of middle school students $\mathcal{I}$. Let $q_{s}$ be the capacity of the high school $s$. Let $\mathcal{M}$ be the finite set of middle schools. The high school $s$ reserves $q_{s}^{m}$ seats for students from the middle school $m$. We assume that $\sum_{m\in M}q_{s}^{m}<q_{s}$, where $q_{s}^{o}=q_{s}-\sum_{m\in M}q_{s}^{m}$ denotes the number of open seats in high school $s$. The function $\tau:\mathcal{I}\rightarrow\mathcal{M}$ is such that $\tau(i)$ denotes the middle school student $i$ belongs to. Let $\mathcal{I}^{m}=\tau^{-1}(m)$ and $I^{m}\equiv I\cap\mathcal{I}^{m}$.
There are two DA-based mechanisms that allocate open and reserved seats simultaneously. In these mechanisms, each high school employs the choice rules $\mathcal{C}^{SimRO}$ and $\mathcal{C}^{SimORO}$ that we define below.
The Simultaneous-reserve-open Choice Rule $\mathcal{C}^{SimRO}$
This choice rule is used in cities such as Guiyan and Tianjin for high school admissions. For any given set $I\subseteq\mathcal{I}$ of students, the choice rule $\mathcal{C}^{SimRO}$ selects the students as follows:
@@
All remaining seats are assigned as open seats to the highest-ranked students among the remaining until all seats are allocated or all students are admitted.
Note that unfilled reserved seats are allocated as open seats in the second stage.
Proposition 2.
$\mathcal{C}^{SimRO}$ is a GSq choice rule.
The Simultaneous-open-reserve-open Choice Rule $\mathcal{C}^{SimORO}$
In Chinese cities Taiyuan, Zengzhou, and Hefei, among many others, for high school admissions, the DA mechanism with the choice rule $\mathcal{C}^{SimORO}$ is used. For any given set $I\subseteq\mathcal{I}$ of students, the choice rule $\mathcal{C}^{SimORO}$ selects the students in three steps.
@@
The unfilled seats in the second stage are allocated as open seats in the third stage.
Proposition 3.
$\mathcal{C}^{SimORO}$ is a GSq choice rule.
Another version of the DA mechanism used for high school admissions in some Chinese cities differs from the two mechanisms above in that each school is divided into two subschools: open and reserved subschools. For school $s\in\mathcal{S}$, the open subschool $s^{o}$ has capacity $q_{S}^{o}$, and the reserved subschool $S^{r}_{s}$ has the capacity vector $q_{s}^{r}=(q_{S}^{m})_{m\in\mathcal{M}}$. We instead model it in matching with contracts framework. Students are required to submit strict preferences over subschools. Then, the student-proposing DA coupled with the following choice rules is run to find a matching.161616pathak2024fair is the first paper to create a hypothetical matching market between individuals and reserve categories in the context of rationing life-saving medical resources. They define a DA-induced matching as the outcome of the individual-proposing DA for a given ”artificial” strict preferences of individuals over categories. In each round of the DA, each unassigned student applies to the highest-ranked subschool in their preferences.
@@
Consider only the contracts in $X_{r}$. For each $m\in\mathcal{M}$, choose contracts of the highest-ranked students such that $(X_{r})_{i}\in I^{m}$ up to capacity $q_{s}^{m}$.
This rule does not allocate reserved positions that are left unfilled. Thus, it is wasteful.
Proposition 4.
$\mathcal{C}^{SimSep}$ is a GSq choice rule.
@@
Consider only the contracts in $X_{r}$. For each $m\in\mathcal{M}$, choose contracts of the highest-ranked students such that $(X_{r})_{i}\in I^{m}$ up to capacity $q_{s}^{m}$.
@@
Proposition 5.
$\mathcal{C}^{SimFlex}$ is a GSq choice rule.
@@
Let $I$ be the set of students and $s$ be the only school. Define $X=I\times\{s\}\times\{o,r\}$. Assume that each student $i\in I$ ranks the contract with the open seat over the contract with the reserved seat. Run the student proposing DA while the school $s$ selects students using $\mathcal{C}^{SimFlex}$. The result of the DA procedure defines $\mathcal{C}^{SimOR}(I)$.
hu2025verifiable show that $\mathcal{C}^{SimOR}$ does not belong to the class of slot-specific priorities choice rules of kominers2016matching.
Proposition 6.
$\mathcal{C}^{SimOR}$ is a GSq choice rule.